OpenStudy (anonymous):

write the complex number in the form a+bi 8(cos30+isin30) I dont get how to apply this to the unit circle

OpenStudy (anonymous):

8.cos(30)=a and 8sin30=b

OpenStudy (anonymous):

okay but the answers arent in degrees

OpenStudy (anonymous):

\[z=8.e^{j.30}\]

OpenStudy (anonymous):

cos(30)=sqrt(3)/2 and sin(30)=1/2

OpenStudy (anonymous):

z=4sqrt(3)+4i

OpenStudy (anonymous):

okay hold on can you explain what you just did

OpenStudy (anonymous):

look to the first thing i told you,i just plug in the values of sin(30) and cos(30)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

.5 and .87

OpenStudy (anonymous):

yes,but you dont need a calc to know this values

OpenStudy (anonymous):

30 is a notable angle

OpenStudy (anonymous):

okay thats where it gets me how do i tell

OpenStudy (anonymous):

oh when i look i see sqrt3/2,1/2

OpenStudy (anonymous):

you need to know the sin,cos and tg of (30,45,60)

OpenStudy (anonymous):

how do i determine that?

OpenStudy (anonymous):

oh thank you thats much appreciated!

OpenStudy (anonymous):

thats a 45 right or 215?

OpenStudy (anonymous):

45

OpenStudy (anonymous):

okay so now hold on let me see if i got this

OpenStudy (anonymous):